# Multilinear Wavelet Compact T(1) Theorem

### v2

@@ 1. Introduction
where $\Lambda^{*,j}$ permutes the 00th and $j$th argument (see (1.3) below). When $m=1$, this is the celebrated $T(1)$ theorem of David and Journé [david84] which was extended to $m\geq 2$ by Grafakos and Torres [GT02] (see also the earlier works [christ1987polynomial, coifman-meyer]) and more recently developed in, for example, [DPLiMV2, DPLiMV1, LN19, li2019bilinear, li2014sharp].
@@
In the linear case, Villaroya in [Vil15] gave a complete characterization of compact Calderón-Zygmund operators on $L^{2}(\mathbb{R}^{d})$, which was further developed in [OV, PPV, SVW, Vil19]. Recently, Mitkovski and Stockdale in [MS23] gave a simplified formulation of the $T(1)$ theorem for compactness of Villaroya. More precisely, they showed that a CZO $T$ is compact if and only if $\mathfrak{b}_{0}$ and $\mathfrak{b}_{1}$ both belong to $\mathrm{CMO}(\mathbb{R}^{d})$ and a vanishing version of the weak boundedness property, called the weak compactness property, is satisfied. Even more recently, Bényi, Li, Oh, and Torres [BLOT24b] have characterized compact CZOs in terms of $L^{2}$ and $L^{\infty}$ testing conditions, in the spirit of Stein’s formulation of the $T(1)$-theorem [stein-book]. These $T(1)$ theorems for compactness are complemented by a compact Rubio De Francia theory of extrapolation due to Hytönen and Lappas in [HL2, HL1], and subsequently developed by others in [COY, CST]. Particularly, in [COY], Cao, Olivo, and Yabuta extended the bilinear results of [HL2] to the multilinear setting and to the quasi-Banach range, in which case the target space can be $L^{r}$ with $r>\frac{1}{m}$. However, due to the difficulties of extrapolating to the upper endpoint in the multilinear setting [li2021endpoint, nieraeth2019quantitative], the results of [COY, HL2] do not consider the case where one (or more) input spaces is $L^{\infty}(\mathbb{R}^{d})$. We point out that our results below do yield compactness when an input space is $L^{\infty}(\mathbb{R}^{d})$.
@@
Our main result is Theorem 2 below which builds on the literature discussed above in the following two noteworthy manners. First, it provides a $T(1)$-type characterization of compact multilinear CZOs. Though extrapolation of compactness had been developed in the multilinear setting as mentioned above, a $T(1)$-type characterization in the vein of Villaroya [Vil15] or the simplified formulation of Mitkovski and Stockdale [MS23] was heretofore unknown. Our characterization is indeed in the simplified form of [MS23] (not relying on compact CZ kernel estimates) while at the same time provides a stronger result than both [BLOT24b] and [MS23] when $m=1$.
@@
Second, we provide a representation formula for all compact CZOs. This is one answer to the search for examples of compact CZOs. Frankly, from Theorem 2, we now know all the examples of $m$-linear compact CZOs. This contribution is new even in the linear case. In words, we can describe our wavelet representation formula in terms of two building blocks whose adjoint operators are well-known examples of compact CZOs. The first class of buiding blocks consists of what we call compact wavelet forms which are almost-orthogonal perturbations of compact approximations of the wavelet resolution of the identity. The second class of building blocks consists of paraproduct forms with $\mathrm{CMO}(\mathbb{R}^{d})$ symbol. The wavelet representation theorem we will prove says that all compact CZOs are finite linear combinations of these two building blocks. We refer to Theorem 2 and Definitions 2.6 and 5.1 below for precise statements.
@@
Moreover, let $\mathbb{B}(z)$ denote a hyperbolic ball in the upper half space of radius $1$ centered at $z=(w,t)$ (see (2.3) below for the precise definition). Fixing three functions $\phi,\psi^{1},\psi^{2}$, define
@@ 2.1. Analysis of the parameter space
Introduce the parameter space
@@
whose elements $z=(w,t)$ act on functions $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{d})$ functions by the formula
@@
For $\mathbf{z}=(\mathbf{w},s)\in\mathbb{Z}^{md}$ and $\zeta=(w_{0},t)\in\mathbb{Z}^{d}$, define for any $0<\delta\leq 1$,
(2.2) $[\mathbf{z},\zeta]_{\delta}=\frac{\min\{s,t\}^{\delta}}{\max\{t,s,|\mathbf{w}-w_{0}|_{2}\}^{md+\delta}}.$
We will also say $\mathbf{z}\geq\zeta$ if $s\geq t$. Notice that if $\delta\geq\delta^{\prime}$ then $[\mathbf{z},\zeta]_{\delta}\leq[\mathbf{z},\zeta]_{\delta^{\prime}}.$ For $M\geq 1$ and $\zeta\in Z^{d}$, introduce $\mathbb{B}_{M}^{m}(\zeta)$ which are the following approximate balls in the hyperbolic metric,
@@ 2.2. Wavelet classes and forms
For $z=(w,t)\in\mathbb{Z}^{d}$, the wavelet class $\Psi^{m,\delta}_{z}$ is defined by
@@
Furthermore, there exists $C_{\infty}>0$ and $\bar{T}$ an extension of $T$ such that
@@ 3. Wavelet representation of compact Calderón-Zygmund forms
If $T$ is a continuous $m$-linear operator from ×$\times_{j=1}^{m}{\mathcal{S}}\to{\mathcal{S}}^{\prime}$ such that $\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle=\Lambda(f_{0},\ldots,f_{m})$ for some CZ form $\Lambda$, then $T$ is said to be a Calderón-Zygmund operator (CZO). Furthermore, let $\phi$ be a mother wavelet and $\psi^{2},\psi^{4}$ be the Schwartz functions from Lemma 2.4. For $j=1,\ldots,m$ define
@@ A CZ form $\Lambda$ is said to be a compact CZ form if for some (all) $M\geq 1$
By the symmetry in Definitions 3.1 and 3.2 above, if $\Lambda$ is a CZ form or compact CZ form, then so is $\Lambda^{\sigma}$ for any $\sigma\in\mathrm{S}_{m+1}$, defined by (1.5). Therefore, if $T$ is a CZO, so is $T^{*,j}$ for any $j=1,\ldots,m$. It is not too difficult to see that if $\mathsf{W}_{\Lambda}^{M}\to 0$ for some $M$ then the same holds for each $M$.
@@
Let $1\leq q_{-}<r<q_{+}\leq\infty$ and $\{f_{z}\in L^{q_{-}}(\mathbb{R}^{d})\cap L^{q_{+}}(\mathbb{R}^{d}):z\in Z^{d}\}$ satisfy
@@
    for $|x-w|\gtrsim t$, $\left|f_{z}(x)\right|\lesssim t^{-\frac{d}{r}}\chi_{z}(x)$.
@@
(iii) ensures that the first quantity is smaller when $z=(w,t)$, $t\sim 1$, and $|w|$ is much larger than $R$. (ii) controls the latter two quantities when $t$ is small or large. ∎
@@
Second, if (3.6) holds, then each $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$ for $j=0,\ldots,m$, and for each $M\geq 1$,
@@
To prove the first conclusion, set $b_{z}^{j}=t^{md-\frac{d}{2}}T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})$. Since $T^{*,j}$ is bounded from ×$\times_{j=1}^{m}L^{qm}(\mathbb{R}^{d})\to L^{q}(\mathbb{R}^{d})$ for any $1<q<\infty$ (being a CZO) we have
@@
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{2m})}\varphi_{z}^{i}&i=1,\ldots,m\\ b_{z}^{0}&i=0;\end{array}\right.\quad p_{i}=2m,\ i=1,\ldots,m;\quad r=2.$
@@
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{4(m-1)})}\varphi_{z}^{i}&i\not\in\{0,j\}\\ t^{d(1-\frac{1}{4})}\varphi_{z}^{j}&i=0\\ b_{z}^{j}&i=j;\end{array}\right.\quad p_{i}=\left\{\begin{array}[]{ll}4(m-1)&i\not\in\{0,j\}\\ 2&i=j;\end{array}\right.\quad r=4$
In either case, as defined, $\{(f^{1}_{z},\ldots,f^{m}_{z}),z\in Z^{d}\}$ is a bounded set in $L^{\vec{p}}$. Crucially, by Lemma 3.7, $f^{0}_{z}$ converges weakly to zero in $L^{r^{\prime}}(\mathbb{R}^{d})$. In the case $f^{0}_{z}=t^{d(1-\frac{1}{r})}\varphi^{j}_{z}$ the conditions are readily verified for $q_{-}=1$ and $q_{+}=\infty$. The more challenging case, when $f^{0}_{z}=b_{z}^{0}$ (the case $j=0$ and $r=2$), also satisfies the assumptions of Lemma 3.7. Indeed, (i) and (ii) are consequences of (3.8) and (iii) can be verified using the kernel estimate (3.1). Now suppose (3.6) fails. Then, by (3.9), there exists $\varepsilon>0$ and $z_{n}\to\infty$ such that
@@
To prove the second statement, Cauchy-Schwarz immediately tells us that (3.7) follows from (3.6). The proof of $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$ is more delicate. To do so, we define the function $\mathfrak{c}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ by $\mathfrak{c}_{j}=T^{*,j}[1,1,\ldots,1]$, where $T[\cdot]$ denotes the now classical extension of $T$ to elements of ×$\times_{j=1}^{m}L^{\infty}(\mathbb{R}^{d})$,
(3.11) $T[f_{1},\ldots,f_{m}]:=\lim_{t\to 0}T(f_{1}\theta_{t},\ldots,f_{m}\theta_{t})-c_{t}$
for suitable constants $c_{t}$ (depending $(m+1)$-linearly on $T,f_{1},\ldots,f_{m}$) with $\lim_{t\to 0}c_{t}=c_{0}\in\mathbb{C}$; see e.g. [GT02]*Lemma 1. Notice that when $f_{j}\in L^{\infty}(\mathbb{R}^{d})\cap\mathcal{S}(\mathbb{R}^{d})$, the definition (3.11) differs from $T(f_{1},\ldots,f_{m})$ but only by a constant. Therefore, when we view (3.11) as an element of $\mathrm{BMO}(\mathbb{R}^{d})$, which consists of equivalence classes modulo constants, there is no contradiction. Clearly $\mathfrak{c}_{j}=\mathfrak{b}_{j}$ in our distributional sense (3.3) as well. It is therefore enough to show $\mathfrak{c}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, which is well-known [Uch78] to be equivalent to showing for each $\varepsilon>0$ there exists $M>0$ such that
@@
Let $\varphi\in L^{\infty}$ satisfy $1_{\{|x|\leq 1\}}\leq\varphi\leq 1_{\{|x|\leq 2\}}$ and, in the sense of (3.11), for any $\zeta\in Z^{d}$, we split $\mathfrak{c}_{j}=E+T[\varphi_{\zeta},\ldots,\varphi_{\zeta}]$, where
$E=T[1-\varphi_{\zeta},1,\ldots,1]+T[\varphi_{\zeta},1-\varphi_{\zeta},1,\ldots,1]+\cdots+T[\varphi_{\zeta},\ldots,\varphi_{\zeta},1-\varphi_{\zeta}].$
Using the kernel estimate (3.2), one can see that for $z=(w,t)$ and $\zeta=(w,at)$, $\mathsf{osc}_{z}(E)\lesssim a^{-\delta}$ for $a\geq 2$. Second, by Hölder’s inequality and the fact that $T$ and $T[\cdot]$ only differ by a constant,
$\mathsf{osc}_{z}(T[\varphi_{\zeta},\ldots,\varphi_{\zeta}])=\mathsf{osc}_{z}(b^{j}_{\zeta})\lesssim_{a}\left\|b^{j}_{\zeta}\right\|_{L^{2}}.$
Therefore, given $\varepsilon>0$, first pick $a$ large enough that $\mathsf{osc}_{z}(E)<\frac{\varepsilon}{2}$. Then, by (3.6), we can pick $M$ large enough that for $\zeta\not\in\mathbb{B}_{M}$, $\mathsf{osc}_{z}(T[\varphi_{\zeta},\ldots,\varphi_{\zeta}])<\frac{\varepsilon}{2}$. But there exists $N\gtrsim_{a}M$ such that if $z=(w,t)\not\in\mathbb{B}_{N}$ then $\zeta=(w,at)\not\in\mathbb{B}_{M}$. ∎
@@
where the sum is taken over all $\sigma\in\mathrm{S}_{m+1}$ and over the combinations $\vec{\psi}_{\mathrm{e}}$ is of the form (3.4) for some $j=1,\ldots,m$ and $\vec{\psi}_{\mathrm{o}}$ is of the same form but with $\psi^{2}$ and $\psi^{4}$ replaced by $\psi^{1}$ and $\psi^{3}$ from Lemma 2.4. Each of these summands (of which there are only finitely many depending on $m$), will now be converted to a compact wavelet form by Proposition 3.10 and Lemma 2.5. Fix now one $\sigma$, $\vec{\psi}_{\mathrm{e}}$, and $\vec{\psi}_{\mathrm{o}}$. We define
[…]

### v3

@@ 1. Introduction
where $\Lambda^{*,j}$ permutes the 0th and $j$th argument (see (1.3) below). When $m=1$, this is the celebrated $T(1)$ theorem of David and Journé [david84] which was extended to $m\geq 2$ by Grafakos and Torres [GT02] (see also the earlier works [christ1987polynomial, coifman-meyer]) and more recently developed in, for example, [DPLiMV2, DPLiMV1, LN19, li2019bilinear, li2014sharp].
@@
Compactness of singular integral and related operators has a long history, with contributions from several distinct directions in harmonic analysis and operator theory. In particular, the study of commutators of Calderón-Zygmund operators, initiated by Uchiyama in the context of Hankel operators [Uch78], has developed into a rich literature; see, e.g., [MR1397362, MR924766, MR1021138] for some foundational work on compactness of commutators and tools used and [BT13, LL22, MR3834654] for some more recent contributions to the area. In another direction, Cordes’ work [Cor75] on compactness of pseudodifferential operators, which lie outside the class of Calderón-Zygmund operators, provides sufficient conditions for compactness in terms of the symbol of the pseudodifferential operator. That line of investigation has seen further recent developments [CIRXY24, CST, TX20]. These results highlight the importance of compactness phenomena in harmonic analysis, though they concern somewhat different mechanisms than those addressed here. In particular, these lines of inquiry are different from what we explore in this paper, as the behavior in question is connected to a particular symbol (function theory), whereas ours is some property intrinsic to a specific operator (its compactness).
@@
In the linear case, Villarroya in [Vil15] gave a complete characterization of compact Calderón-Zygmund operators on $L^{2}(\mathbb{R}^{d})$, which was further developed in [OV, PPV, SVW, Vil19]. Recently, Mitkovski and Stockdale in [MS23] gave a simplified formulation of the $T(1)$ theorem for compactness of Villarroya. More precisely, they showed that a CZO $T$ is compact if and only if $\mathfrak{b}_{0}$ and $\mathfrak{b}_{1}$ both belong to $\mathrm{CMO}(\mathbb{R}^{d})$ and a vanishing version of the weak boundedness property, called the weak compactness property, is satisfied. Even more recently, Bényi, Li, Oh, and Torres [BLOT24b] have characterized compact CZOs in terms of $L^{2}$ and $L^{\infty}$ testing conditions, in the spirit of Stein’s formulation of the $T(1)$-theorem [stein-book]. These $T(1)$ theorems for compactness are complemented by a compact Rubio De Francia theory of extrapolation due to Hytönen and Lappas in [HL2, HL1], and subsequently developed by others in [COY, CST]. Particularly, in [COY], Cao, Olivo, and Yabuta extended the bilinear results of [HL2] to the multilinear setting and to the quasi-Banach range, in which case the target space can be $L^{r}$ with $r>\frac{1}{m}$. However, due to the difficulties of extrapolating to the upper endpoint in the multilinear setting [li2021endpoint, nieraeth2019quantitative], the results of [COY, HL2] do not consider the case where one (or more) input spaces is $L^{\infty}(\mathbb{R}^{d})$. We point out that our results below do yield compactness when an input space is $L^{\infty}(\mathbb{R}^{d})$.
@@
Our main result is Theorem 2 below which builds on the literature discussed above in the following two noteworthy manners. First, it provides a $T(1)$-type characterization of compact multilinear CZOs. Though extrapolation of compactness had been developed in the multilinear setting as mentioned above, a $T(1)$-type characterization in the vein of Villarroya [Vil15] or the simplified formulation of Mitkovski and Stockdale [MS23] was heretofore unknown. Our characterization is indeed in the simplified form of [MS23] (not relying on compact CZ kernel estimates) while at the same time provides a stronger result than both [BLOT24b] and [MS23] when $m=1$. In particular, the implication B. to E. in Theorems 2 and 3 resolves the question posed in [BLOT24b]*Remark 3.6, while at the same time extends the $L^{2}$-testing theorem to the multilinear setting as well.
@@
Second, we provide a representation formula for all compact CZOs. This is one answer to the search for examples of compact CZOs. Frankly, from Theorem 2, we now know all the examples of $m$-linear compact CZOs. This contribution is new even in the linear case. In words, we can describe our wavelet representation formula in terms of two building blocks whose adjoint operators are well-known examples of compact CZOs. The first class of building blocks consists of what we call compact wavelet forms which are almost-orthogonal perturbations of compact approximations of the wavelet resolution of the identity. The second class of building blocks consists of paraproduct forms with $\mathrm{CMO}(\mathbb{R}^{d})$ symbol. The wavelet representation theorem we will prove says that all compact CZOs are finite linear combinations of these two building blocks. We refer to Theorem 2 and Definitions 2.6 and 5.1 below for precise statements.
@@
Moreover, let $\mathbb{B}(z)$ denote a hyperbolic ball in the upper half space of radius $1$ centered at $z=(w,t)$ (see (2.3) and (2.4) below for the precise definitions). Fixing three functions $\phi,\psi^{1},\psi^{2}$, define
@@ 2.1. Analysis of the parameter space
For a positive integer $D$, introduce the parameter space
@@
whose elements $z=(w,t)$ act on functions $f\in L^{1}_{\mathrm{loc}}(\mathbb{R}^{D})$ by the formula
@@
For $\mathbf{z}=(\mathbf{w},s)\in Z^{md}$ and $\zeta=(w_{0},t)\in Z^{d}$, define for any $0<\delta\leq 1$,
(2.2) $\mathbf{z},\zeta]_{\delta}=\frac{\min\{s,t\}^{\delta}}{\max\{t,s,|\mathbf{w}-w_{0}|_{2}\}^{md+\delta}}.$
We will also say $\mathbf{z}\geq\zeta$ if $s\geq t$. Notice that if $\delta\geq\delta^{\prime}$ then $\mathbf{z},\zeta]_{\delta}\leq[\mathbf{z},\zeta]_{\delta^{\prime}}.$ For $M\geq 1$ and $\zeta\in Z^{d}$, introduce $\mathbb{B}_{M}^{m}(\zeta)$ which are the following approximate balls in the hyperbolic metric,
@@ 2.2. Wavelet classes and forms
For $z=(w,t)\in Z^{d}$, the wavelet class $\Psi^{m,\delta}_{z}$ is defined by
@@
Furthermore, there exist $C_{\infty}>0$ and $\bar{T}$ an extension of  $T$ such that
@@ 3. Wavelet representation of compact Calderón-Zygmund forms
If $T$ is a continuous $m$-linear operator from $\times_{j=1}^{m}{\mathcal{S}}\to{\mathcal{S}}^{\prime}$ such that $\left\langle f_{0},T(f_{1},\ldots,f_{m})\right\rangle=\Lambda(f_{0},\ldots,f_{m})$ for some CZ form $\Lambda$, then $T$ is said to be a Calderón-Zygmund operator (CZO). Furthermore, let $\phi$ be a mother wavelet and $\psi^{2},\psi^{4}$ be the Schwartz functions from Lemma 2.4. For $j=1,\ldots,m$ define
@@ A CZ form $\Lambda$ is said to be a compact CZ form if for some (all) $M\geq 1$
By the symmetry in Definitions 3.1 and 3.2 above, if $\Lambda$ is a CZ form or compact CZ form, then so is $\Lambda^{\sigma}$ for any $\sigma\in\mathrm{S}_{m+1}$, defined by (1.5). Therefore, if $T$ is a CZO, so is $T^{*,j}$ for any $j=1,\ldots,m$. It is not too difficult to see that if $\mathsf{W}_{\Lambda}^{M}\to 0$ for some $M$ then the same holds for each $M$. Therefore, when convenient, we may reduce to the case $M=1$ and utilize the shorthand $\mathsf{W}_{\Lambda}(\zeta)=\mathsf{W}_{\Lambda}^{1}(\zeta)$.
@@
Let $1\leq q_{-}<r<q_{+}\leq\infty$ and $\{f_{z}\in L^{q_{-}}(\mathbb{R}^{d})\cap L^{q_{+}}(\mathbb{R}^{d}):z\in Z^{d}\}$ be a collection of functions satisfying
@@
    There exists $c>0$ such that for $|x-w|\geq c\cdot t$, $\left|f_{z}(x)\right|\lesssim t^{-\frac{d}{r}}\chi_{z}(x)$.
@@
(iii) ensures that the first quantity is smaller when $z=(w,t)$, $t\sim 1$, and $|w|$ is much larger than R+cR{+c}. (ii) controls the latter two quantities when $t$ is small or large. ∎
@@
Second, if (3.6) holds, then each $\mathfrak{b}_{j}$ belongs to $\mathrm{CMO}(\mathbb{R}^{d})$ for $j=0,\ldots,m$, and for each $M\geq 1$,
@@
To prove the first conclusion, set $b_{z}^{j}=t^{md-\frac{d}{2}}T^{*,j}(\varphi^{1}_{z},\ldots,\varphi^{m}_{z})$. Since $T^{*,j}$ is bounded from $\times_{j=1}^{m}L^{qm}(\mathbb{R}^{d})\to L^{q}(\mathbb{R}^{d})$ for any $1<q<\infty$ (being a CZO) we have
@@
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{2m})}\varphi_{z}^{i},&i=1,\ldots,m;\\ b_{z}^{0},&i=0;\end{array}\right.\quad p_{i}=2m,\ i=1,\ldots,m;\quad r=2.$
@@
$f_{z}^{i}=\left\{\begin{array}[]{ll}t^{d(1-\frac{1}{4(m-1)})}\varphi_{z}^{i},&i\not\in\{0,j\};\\ t^{d(1-\frac{1}{4})}\varphi_{z}^{j},&i=0;\\ b_{z}^{j},&i=j;\end{array}\right.\quad p_{i}=\left\{\begin{array}[]{ll}4(m-1),&i\not\in\{0,j\};\\ 2,&i=j;\end{array}\right.\quad r=4.$
In either case, as defined, $\{(f^{1}_{z},\ldots,f^{m}_{z}),z\in Z^{d}\}$ is a bounded set in $L^{\vec{p}}$. Crucially, by Lemma 3.7, $f^{0}_{z}$ converges weakly to zero in $L^{r^{\prime}}(\mathbb{R}^{d})$. In the case $f^{0}_{z}=t^{d(1-\frac{1}{r})}\varphi^{j}_{z}$ the conditions are readily verified for $q_{-}=1$ and $q_{+}=\infty$. The more challenging case, when $f^{0}_{z}=b_{z}^{0}$ (the case $j=0$ and $r=2$), also satisfies the assumptions of Lemma 3.7. Indeed, (i) and (ii) are consequences of (3.8). To establish (iii), take $c>0$ large enough that all $\varphi^{j}$ are supported in the ball $B(0,\frac{c}{2})$. Then, since $\varphi^{j}_{z}$ is supported in the ball $B(w,\frac{ct}{2})$, for $x$ outside $B(w,ct)$, (3.1) implies $K(x,\mathbf{x})\lesssim|x-w|^{-md}$. The condition (iii) is then a consequence of
$b^{0}_{z}(x)=t^{md-\frac{d}{2}}\int_{\mathbb{R}^{md}}K(x,\mathbf{x})\prod_{j=1}^{m}\varphi_{z}^{j}(x_{j})\,{\rm d}\mathbf{x}\lesssim t^{-\frac{d}{2}}\left(\frac{|x-w|}{t}\right)^{-md}\lesssim t^{-\frac{d}{2}}\chi_{z}(x).$
Now suppose (3.6) fails. Then, by (3.9), there exist $\varepsilon>0$ and $z_{n}\to\infty$ such that
@@
To prove the second statement, Cauchy-Schwarz immediately tells us that (3.7) follows from (3.6). The proof of $\mathfrak{b}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$ is more delicate. To do so, we define the function $\mathfrak{c}_{j}\in\mathrm{BMO}(\mathbb{R}^{d})$ by $\mathfrak{c}_{j}=T^{*,j}[1,1,\ldots,1]$, where $T[\cdot]$ denotes the now classical extension of $T$ to elements of $\times_{j=1}^{m}L^{\infty}(\mathbb{R}^{d})$,
(3.11) $T[f_{1},\ldots,f_{m}]:=\lim_{t\to 0}T(f_{1}\theta_{t},\ldots,f_{m}\theta_{t})-c_{t},$
where the limit exists, distributionally, locally in $L^{1}$, and pointwise a.e., and $c_{t}$ are suitable constants depending $(m+1)$-linearly on $T,f_{1},\ldots,f_{m}$ with $\lim_{t\to 0}c_{t}=c_{0}\in\mathbb{C}$; see e.g. [GT02]*Lemma 1 and Proposition 1. Notice that when $f_{j}\in L^{\infty}(\mathbb{R}^{d})\cap\mathcal{S}(\mathbb{R}^{d})$, the definition (3.11) differs from $T(f_{1},\ldots,f_{m})$ but only by a constant. Therefore, when we view (3.11) as an element of $\mathrm{BMO}(\mathbb{R}^{d})$, which consists of equivalence classes modulo constants, there is no contradiction. Clearly $\mathfrak{c}_{j}=\mathfrak{b}_{j}$ in our distributional sense (3.3) as well. It is therefore enough to show $\mathfrak{c}_{j}\in\mathrm{CMO}(\mathbb{R}^{d})$, which is well-known [Uch78] to be equivalent to showing for each $\varepsilon>0$ there exists $M>0$ such that
@@
Let $\varphi\in L^{\infty}$ satisfy $1_{\{|x|\leq 1\}}\leq\varphi\leq 1_{\{|x|\leq 2\}}$ and for $\zeta=(\xi,\sigma)\in Z^{d}$ to be chosen later, we set $\varphi_{\zeta}=\sigma^{d}{\mathsf{Sy}}_{\zeta}\varphi$. In the sense of (3.11), split $\mathfrak{c}_{j}=E+T^{*,j}[\varphi_{\zeta},\ldots,\varphi_{\zeta}]$, where
$E=T^{*,j}[1-\varphi_{\zeta},1,\ldots,1]+T^{*,j}[\varphi_{\zeta},1-\varphi_{\zeta},1,\ldots,1]+\cdots+T^{*,j}[\varphi_{\zeta},\ldots,\varphi_{\zeta},1-\varphi_{\zeta}].$
Let now $z=(w,t)$ and we will estimate $\mathsf{osc}_{z}(E)$ using the smoothness estimate for the kernel (3.2). To this end, for $s>0$, let $E_{s}^{k}=T^{*,j}(\varphi_{\zeta}\theta_{s},\ldots,\varphi_{\zeta}\theta_{s},(1-\varphi_{\zeta})\theta_{s},\theta_{s},\ldots,\theta_{s})$, where $(1-\varphi_{\zeta})$ is in the $k$-th position. Then, for any $a\geq 2$, set $\zeta=(w,at)$. The following estimate holds uniformly over $s>0$, $a\geq 2$, and $x,y\in B(w,t)$,
$|E_{s}^{k}(x)-E_{s}^{k}(y)|\lesssim t^{\delta}\int_{\begin{subarray}{c}\mathbf{x}\in\mathbb{R}^{md}\\ |x_{k}-x|\gtrsim at\end{subarray}}|x-\mathbf{x}|_{2}^{-md-\delta}\,{\rm d}\mathbf{x}.$
The remaining integral is controlled by $(at)^{-\delta}$. Passing to the limit, $|E(x)-E(y)|\lesssim a^{-\delta}$ and hence $\mathsf{osc}_{z}(E)\lesssim a^{-\delta}$. Second, by Hölder’s inequality and the fact that $T$ and $T[\cdot]$ only differ by a constant,
$\mathsf{osc}_{z}(T^{*,j}[\varphi_{\zeta},\ldots,\varphi_{\zeta}])=(at)^{\frac{d}{2}}\mathsf{osc}_{z}(b^{j}_{\zeta})\lesssim_{a}\left\|b^{j}_{\zeta}\right\|_{L^{2}}.$
Therefore, given $\varepsilon>0$, first pick $a$ large enough that $\mathsf{osc}_{z}(E)<\frac{\varepsilon}{2}$. Then, by (3.6), we can pick $M$ large enough that for $\zeta\not\in\mathbb{B}_{M}$, $\mathsf{osc}_{z}(T^{*,j}[\varphi_{\zeta},\ldots,\varphi_{\zeta}])<\frac{\varepsilon}{2}$. But there exists $N\gtrsim_{a}M$ such that if $z=(w,t)\not\in\mathbb{B}_{N}$ then $\zeta=(w,at)\not\in\mathbb{B}_{M}$. ∎
@@
where the sum is taken over all $\sigma\in\mathrm{S}_{m+1}$ and over the combinations $\vec{\psi}_{\mathrm{e}}$ of the form (3.4) for some $j=1,\ldots,m$ and $\vec{\psi}_{\mathrm{o}}$ is of the same form but with $\psi^{2}$ and $\psi^{4}$ replaced by $\psi^{1}$ and $\psi^{3}$ from Lemma 2.4. Each of these summands (of which there are only finitely many depending on $m$), will now be converted to a compact wavelet form by Proposition 3.10 and Lemma 2.5. Fix now one $\sigma$, $\vec{\psi}_{\mathrm{e}}$, and $\vec{\psi}_{\mathrm{o}}$. We define
$\vartheta_{\zeta}=\int_{\begin{subarray}{c}\mathbf{z}\in Z^{md}\\ \mathbf{z}\geq\zeta\end{subarray}}\Lambda^{\sigma}((\vec{\psi}_{\mathrm{e}})_{\mathbf{z}},\phi_{\zeta})(\vec{\psi}_{\mathrm{o}})_{\mathbf{z}}\,d\mu(\mathbf{z})$